# Virne: A Comprehensive Benchmark for Deep RL-based Network Resource Allocation in NFV

### v1

@@ Abstract
Resource allocation (RA) is critical to efficient service deployment in Network Function Virtualization (NFV), a transformative networking paradigm. Recently, deep Reinforcement Learning (RL)-based methods have been showing promising potential to address this complexity. However, the lack of a systematic benchmarking framework and thorough analysis hinders the exploration of emerging networks and the development of more robust algorithms while causing inconsistent evaluation. In this paper, we introduce Virne, a comprehensive benchmarking framework for the NFV-RA problem, with a focus on supporting deep RL-based methods. Virne provides customizable simulations for diverse network scenarios, including cloud, edge, and 5G environments. It also features a modular and extensible implementation pipeline that supports over 30 methods of various types, and includes practical evaluation perspectives beyond effectiveness, such as solvability, generalization, and scalability. Furthermore, we conduct in-depth analysis through extensive experiments to provide valuable insights into performance trade-offs for efficient implementation and offer actionable guidance for future research directions. Overall, with its diverse simulations, rich implementations, and extensive evaluation capabilities, Virne could serve as a comprehensive benchmark for advancing NFV-RA methods and deep RL applications. The code is publicly available at https://github.com/GeminiLight/virne.
@@ 1 Introduction
Network Function Virtualization (NFV) has emerged as an essential enabler for modern networks, such as cloud data centers, edge computing and 5G, due to its remarkable flexibility and scalability vne-cn-2018-survey-nfv . By transforming traditional hardware-bound network service functions into flexible software modules, NFV enables the deployment of Virtual Network Functions (VNFs) on general-purpose servers vne-pieee-2020-survey-nfv-sdn . A central challenge in NFV is Resource Allocation (NFV-RA) that is essential for effective resource management and ensuring service quality vne-tpds-2021-survey-nfv-ra . As illustrated in Figure 1, NFV-RA involves mapping service requests (modeled as virtual networks of interconnected VNFs) onto shared physical infrastructure while satisfying various constraints. It is an NP-hard combinatorial optimization problem vne-ton-2020-nphard , highlighting the need for efficient solution strategies.
Traditional approaches to NFV-RA, such as exact solvers vne-infocom-2009-rd ; vne-jsac-2018-ilp that seek optimal solutions or heuristics vne-infocom-2014-grc ; vne-iotj-2018-nrm ; vne-tpds-2023-nea dependent on manual design, suffer from neither excessive time consumption nor suboptimal performance. Recently, Reinforcement Learning (RL) has shown promise in solving NFV-RA, which enables the autonomous learning of efficient heuristics by interacting with simulated network environments vne-tcyb-2017-mcts ; vne-iwqos-2019-pg-mlp ; vne-tnsm-2020-pg-rnn ; vne-icc-2021-drl-sfcp ; vne-jsac-2020-a3c-gcn ; vne-tnsm-2022-pg-cnn ; vne-infocom-2019-dcnn ; vne-tsc-2023-hrl-acra ; kdd-2023-gal-vne ; vne-ijcai-2024-flag-vne ; vne-comst-2024-ai-vne . This training paradigm operates without requiring high-quality labeled datasets, which are difficult to obtain due to the computational complexity and privacy concerns, making it particularly well-suited for NFV-RA.
@@
In this paper, we introduce Virne, a comprehensive benchmarking framework for NFV-RA that offers diverse simulations, unified implementations and thorough evaluations. Firstly, Virne is designed to serve as a unified and readily accessible framework for researchers from both the machine learning and networking communities. It provides a highly customizable simulation environment capable of accurately modeling a wide array of NFV scenarios, from cloud data centers to edge and 5G networks, allowing for the exploration of various resource types, constraints, and service requirements. Secondly, Virne features a modular architecture that simplifies the implementation of various NFV-RA algorithms, which facilitates both efficient utilization and new algorithm development. It includes over 30 NFV-RA algorithms, covering both exact, heuristic and advanced learning-based methods. Beyond standard performance metrics, Virne enables in-depth analysis through practical evaluation perspectives such as solution feasibility, generalization across varying network conditions, and scalability with increasing problem size. Through extensive empirical studies conducted within Virne, we offer valuable insights into the effectiveness and characteristics of various algorithms, providing data-driven guidance for future research directions and practical deployments.
@@
  • Comprehensive Simulation and Gym-style Environments. We propose Virne, the most comprehensive benchmark for NFV-RA to date, along with gym-style environments. It offers highly customizable simulation and thorough evaluation across diverse scenarios.
  • Unified Implementations of Deep RL-based Methods. To facilitate convenient utilization and future research, we design a modular and easily expandable pipeline that streamlines the implementation process and integrates a wide range of NFV-RA algorithms.
  • Insightful Findings on ML Applications. Through extensive experiments from multiple perspectives, our empirical results reveal the impact of different modules and nuanced performance comparisons, providing actionable insights for future applied ML research.
@@
VNE-Sim simtools-2016-vne-sim Cloud 3 heuristics Effectiveness 2014
ALEVIN vne-netwks-2014-alevin2 Cloud 5 heuristics Effectiveness 2016
ALib vne-sigccr-2019-alib-lp Cloud 1 exact Effectiveness 2019
SFCSim vne-cc-2022-sfc-smi Cloud 3 heuristics Effectiveness 2022
Iflye vne-2021-iflye Cloud 3 heuristics Effectiveness 2024
Virne (Ours) Cloud, Edge, 5G Slicing, etc. 30+ algorithms (10+ non-RL) Effectiveness & 3 Practicality 2025
@@ 2.1 Basic Definition for Cost Optimization
System Model
In NFV, user network services and the physical infrastructure are virtualized as virtual networks (VNs) and physical network (PN), respectively. As illustrated in Figure 1, in online network systems, service requests represented as VNs are continuously arriving at PN, seeking the physical resource with specific Quality-of-Service requirements. Each arrived VN and the corresponding snapshot of PN consist of an instance $I=(\mathcal{G}_{v},\mathcal{G}_{p})$, where the PN $\mathcal{G}_{p}$ and VN $\mathcal{G}_{v}$ are modeled as undirected graphs, $\mathcal{G}_{p}=(\mathcal{N}_{p},\mathcal{L}_{p})$ and $\mathcal{G}_{v}=(\mathcal{N}_{v},\mathcal{L}_{v},\omega,\varpi)$, respectively. Here, $\mathcal{N}_{p}$ and $\mathcal{L}_{p}$ denote the sets of physical nodes and links, indicating servers and their interconnections; $\mathcal{N}_{v}$ and $\mathcal{L}_{v}$ denote the sets of virtual nodes and links, representing services and their relationships; $\omega$ and $\varpi$ denote the arrival time and lifetime of VN. We denote $C(n_{p})$ as available computing resources for physical node $n_{p}\in\mathcal{N}_{p}$, and $B(l_{p})$ as available bandwidth resources of physical link $l_{p}\in\mathcal{L}_{p}$. Besides, $C(n_{v})$ and $B(l_{v})$ denote the demands for computing resource by a virtual node $n_{v}\in\mathcal{N}_{v}$ and bandwidth resource by a virtual link $l_{v}\in\mathcal{L}_{v}$. Over a specific period, we collect all instances into a set $\mathcal{I}$.
Embedding Constraints
Mapping a VN $\mathcal{G}_{v}$ onto the sub-PN $\mathcal{G}_{p}\prime$ is formulated as a mapping function $f_{\mathcal{G}}:\mathcal{G}_{v}\rightarrow\mathcal{G}_{p}\prime$. It consists of two sub-mapping processes, i.e., node mapping $f_{\mathcal{N}}$ and link mapping $f_{\mathcal{L}}$. Node mapping $f_{\mathcal{N}}$, aims to find a suitable physical node $n_{p}=f_{\mathcal{N}}(n_{v})$ to place each virtual node $n_{v}$ while adhering to the one-to-one mapping and resource availability constraints. Concretely, virtual nodes in the same VN must be placed in different physical nodes and each physical node can host at most one virtual node. And the physical node $n_{p}$ must have enough available resources to place the corresponding virtual node $n_{v}$, i.e., $C(n_{v})\leq C(n_{p}),\forall n_{v}\in\mathcal{N}_{v}$. Link mapping aims to find a connected physical path $\rho_{p}=f_{\mathcal{L}}(l_{v})$ to route each virtual link $l_{v}$ while satisfying the path connectivity and resource capacity constraints. Concretely, the physical path $\rho_{p}$ should connect physical nodes hosting the two endpoints of virtual link $l_{v}$. And each physical link $l_{p}\in\rho_{p}$ of physical path $\rho_{p}$ should have enough bandwidth to route the corresponding virtual link $l_{v}$, i.e., $B(l_{v})\leq B(l_{p}),\forall l_{v}\in\mathcal{L}_{v},l_{p}\in\rho_{p}$.
Optimization Objective
Considering the randomness of online service requests, NFV-RA mainly aims to maximize the resource utilization of each instance $I$, which facilitates long-term resource profit and request acceptance vne-tnsm-2022-pg-cnn ; vne-tpds-2023-att ; vne-tsc-2023-gcn-mask . To assess the quality of solution $S=f_{\mathcal{G}}(I)$ to the instance $I$, the revenue-to-cost ratio (R2C) is a widely used metric, defined as follows:
@@
Here, $\varkappa$ is a binary variable indicating the feasibility of the solution $S$: $\varkappa=0$ if $S$ violates some constraints, and $\varkappa=1$ otherwise. $\text{REV}(S)$ and $\text{COST}(S)$ denotes the generated revenue and incurred cost by embedding VN $\mathcal{G}_{v}$. If $\varkappa=1$, $\text{REV}(S)$ denotes the revenue from the VN, calculated as $\sum_{n_{v}\in\mathcal{N}_{v}}{C}(n_{v})+\sum_{l_{v}\in\mathcal{L}_{v}}{B}(l_{v})$ and $\text{COST}(S)$ denotes the resource consumption of PN, calculated as $\sum_{n_{v}\in\mathcal{N}_{v}}{C}(n_{v})+\sum_{l_{v}\in\mathcal{L}_{v}}(|f_{\mathcal{L}}(l_{v})|\times{B}(l_{v}))$. Here, $|f_{\mathcal{L}}(l_{v})|$ quantifies the length of the physical path $\rho_{p}$ routing the virtual link $l_{v}$. See Appendix A.1 for the detailed formulation.
@@ 3 Virne: A Comprehensive NFV-RA Benchmark
To provide high-quality simulation and implementation, we introduce three design principles for Virne, following established software engineering practices van2008software : (a) Versatile customization allows the platform to meet diverse simulation needs on varying network scenarios and conditions, ensuring adaptability. (b) Scalable modularity built platforms with a modular architecture to support flexible configurations and easy extensibility. (c) Intuitive usability prioritizes a user-friendly interface, enabling users to focus on experiment outcomes rather than implementation complexities. Guided by these principles, as illustrated in Figure 2, we implement Virne with five key modules as follows.
@@ 3.1 Simulation Configuration
Modern network systems are diverse and intricate, associated with various realistic factors such as resource availability and service requirements. To support high customization for simulating different network scenarios and conditions, Virne abstracts both the PN and VN as graphs, with customizable attributes for nodes, links, and the overall network. Specifically, Virne provides the following key customizable elements: (a) Network Topologies. Users can select from various topological synthesis methods or real-world physical infrastructure topologies, such as those from SDNLib [^1] https://sndlib.put.poznan.pl/. (b) Resource Availability. Virne enables users to define multiple resource types (e.g., CPU, GPU, bandwidth) and their availability across different levels of the network (i.e., nodes, links and graph), allowing the reflection of the specific resource characteristics of different network environments (c) Service Requirements. Users can specify additional service requirements, such as latency, energy efficiency, and reliability, to ensure that the simulation reflects the needs of different network applications. By adjusting these parameters, Virne provides the flexibility to simulate a wide array of network scenarios, including cloud-based infrastructures, edge computing and 5G networks. This customization enables comprehensive testing, accommodating fluctuating resource demands and evolving network conditions. We provide configuration examples in Figure 4 and Appendix C.1.
@@ 3.3.1 NFV-RA as a Markov Decision Process
To address the randomness of online networks, most existing works model NFV-RA solution construction as a Markov Decision Process (MDP), which sequentially selects a physical node to place a virtual node at each decision step $t$, until all virtual nodes are placed or any constraints are violated.
@@
At each decision timestep $y$, observing the state $s_{t}$ of the environment, the agent takes an action $a_{t}\sim\pi(\cdot|s_{t})$ according to a policy $\pi_{\theta}$ parameterized by $\theta$. Then, the environment will transit to a new state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$ and feedback a reward $R(s_{t},a_{t})$ and. During interactions, a trajectory memory collects these step information as experiences. Using these experiences, the objective of the agent is to learn an optimal policy for resource allocation that maximizes the expected sum of discounted reward, i.e., the solution quality:
$\pi_{\theta}^{*}=\arg\max_{\pi_{\theta}}\mathbb{E}_{\pi_{\theta}}\left[\sum_{i=0}^{T}\gamma^{i}R(s_{t},a_{t})\right]$ (2)
Note that the above MDP introduced is the most adopted version. As a holistic framework, Virne also supports the emerging MDP version in recent studies. Due to the presentation clarity and page limit, we offer other MDP versions of NFV-RA in Appendix C.2.

### v2

# Virne: A Comprehensive Benchmark for RL-based Network Resource Allocation in NFV

@@ Abstract
Resource allocation (RA) is critical to efficient service deployment in Network Function Virtualization (NFV), a transformative networking paradigm. This task is termed NFV-RA. Recently, deep Reinforcement Learning (RL)-based methods have been showing promising potential to address this combinatorial complexity of constrained cross-graph mapping. However, RL-driven NFV-RA research lacks a systematic benchmark for comprehensive simulation and rigorous evaluation. This gap hinders in-depth performance analysis and slows algorithm development for emerging networks, resulting in fragmented assessments. In this paper, we introduce Virne, a comprehensive benchmarking framework designed to accelerate the research and application of deep RL for NFV-RA. Virne provides customizable simulations for diverse network scenarios, including cloud, edge, and 5G environments. It features a modular and extensible implementation pipeline that integrates over 30 methods of various types. Virne also establishes a rigorous evaluation protocol that extends beyond online effectiveness to include practical perspectives such as solvability, generalizability, and scalability. Furthermore, we conduct in-depth analysis through extensive experiments to provide valuable insights into performance trade-offs for efficient implementation and offer actionable guidance for future research directions. Overall, with its capabilities of diverse simulations, rich implementations, and thorough evaluation, Virne could serve as a comprehensive benchmark for advancing NFV-RA methods and deep RL applications. The code and resources are available at https://github.com/GeminiLight/Virne.
@@ 1 Introduction
Network Function Virtualization (NFV) has emerged as an essential enabler for modern networks, such as cloud data centers, edge computing and 5G, due to its remarkable flexibility and scalability (Yi et al., 2018). By transforming traditional hardware-bound network services into flexible software modules, NFV enables the deployment of Virtual Network Functions (VNFs) on general-purpose servers (Zhuang et al., 2020). A central challenge in NFV is Resource Allocation (NFV-RA) that is essential for effective resource management and service quality (Yang et al., 2021). As illustrated in Figure 1, NFV-RA involves mapping service requests (modeled as virtual networks of interconnected VNFs) onto shared physical infrastructure while satisfying constraints. It is an NP-hard combinatorial optimization problem (Rost and Schmid, 2020), highlighting the need for efficient solutions.
Traditional approaches to NFV-RA, such as exact solvers (Chowdhury et al., 2009; Shahriar et al., 2018) that seek optimal solutions or heuristics (Gong et al., 2014; Fan et al., 2023) dependent on manual design, suffer from neither excessive time consumption nor suboptimal performance. Recently, Reinforcement Learning (RL) has shown promise in solving NFV-RA, which enables the autonomous learning of efficient heuristics by interacting with simulated network environments (Haeri and Trajković, 2017; Yan et al., 2020; Wang et al., 2024c; Wu et al., 2024b). This training paradigm operates without requiring high-quality labeled datasets, which are difficult to obtain due to the computational complexity and privacy concerns, making it particularly well-suited for NFV-RA.
@@
In this paper, we introduce Virne, a comprehensive benchmarking framework for NFV-RA that offers diverse simulations, unified implementations and thorough evaluations. Firstly, Virne is designed to serve as a unified and readily accessible framework for researchers from both the machine learning and networking communities. It provides a highly customizable simulation environment capable of accurately modeling a wide array of NFV scenarios, from cloud data centers to edge and 5G networks, allowing for the exploration of various resource types, constraints, and service requirements. Secondly, Virne features a modular architecture that simplifies the implementation of various NFV-RA algorithms, which facilitates both efficient utilization and new algorithm development. It includes over 30 NFV-RA algorithms, covering both exact, heuristic and advanced learning-based methods. Thirdly, beyond standard performance metrics, Virne enables in-depth analysis through practical evaluation perspectives such as solution feasibility, generalization across varying network conditions, and scalability with increasing problem size. Finally, through extensive empirical studies conducted within Virne, we offer valuable insights into the effectiveness and characteristics of various algorithms, providing data-driven guidance for future research directions and practical deployments.
@@
  • Comprehensive Simulations. Virne is the most comprehensive benchmark for NFV-RA to date, along with gym-style environments supporting highly customizable simulations.
  • Streamlined Implementations. To facilitate community use, we design a modular and easily expandable implementation pipeline and integrate a wide range of NFV-RA algorithms.
  • Emerging Evalutions. Beyond online effectiveness alone, Virne enables practical evaluations of solution feasibility, generalization across diverse network conditions, and scalability.
  • Insighful Findings. Extensive results reveal the impact of different modules and nuanced performance comparisons, providing actionable insights for future applied ML research.
@@
VNE-Sim (Haeri and Trajković, 2016) Cloud 3 heuristics Effectiveness 2014
ALEVIN (Beck et al., 2014) Cloud 5 heuristics Effectiveness 2016
ALib (Rost et al., 2019) Cloud 1 exact Effectiveness 2019
SFCSim (Xu et al., 2022) Cloud 3 heuristics Effectiveness 2022
Iflye (Tomaszek, 2021) Cloud 3 heuristics Effectiveness 2024
[…]
@@ 2.1 Basic Definition for Cost Optimization
System Model. In NFV, user network services and the physical infrastructure are virtualized as virtual networks (VNs) and physical network (PN), respectively. As illustrated in Figure 1, in online network systems, service requests represented as VNs are continuously arriving at PN, seeking the physical resource with specific service requirements. Each arrived VN and the corresponding snapshot of PN consist of an instance $I=(\mathcal{G}_{v},\mathcal{G}_{p})$, where the PN $\mathcal{G}_{p}$ and VN $\mathcal{G}_{v}$ are modeled as undirected graphs, $\mathcal{G}_{p}=(\mathcal{N}_{p},\mathcal{L}_{p})$ and $\mathcal{G}_{v}=(\mathcal{N}_{v},\mathcal{L}_{v},\omega,\varpi)$, respectively. Here, $\mathcal{N}_{p}$ and $\mathcal{L}_{p}$ denote the sets of physical nodes and links, indicating servers and their interconnections; $\mathcal{N}_{v}$ and $\mathcal{L}_{v}$ denote the sets of virtual nodes and links, representing services and their relationships; $\omega$ and $\varpi$ denote the arrival time and lifetime of VN. We denote $C(n_{p})$ as available computing resources for physical node $n_{p}\in\mathcal{N}_{p}$, and $B(l_{p})$ as available bandwidth resources of physical link $l_{p}\in\mathcal{L}_{p}$. Besides, $C(n_{v})$ and $B(l_{v})$ denote the demands for computing resource by a virtual node $n_{v}\in\mathcal{N}_{v}$ and bandwidth resource by a virtual link $l_{v}\in\mathcal{L}_{v}$. Over a specific period, we collect all instances into a set $\mathcal{I}$.
@@
Embedding Constraints. Mapping a VN $\mathcal{G}_{v}$ onto the sub-PN $\mathcal{G}_{p}\prime$ is formulated as a mapping function $f_{\mathcal{G}}:\mathcal{G}_{v}\rightarrow\mathcal{G}_{p}\prime$. It consists of two sub-mapping processes, i.e., node mapping $f_{\mathcal{N}}$ and link mapping $f_{\mathcal{L}}$. Node mapping $f_{\mathcal{N}}$, aims to find a suitable physical node $n_{p}=f_{\mathcal{N}}(n_{v})$ to place each virtual node $n_{v}$ while adhering to the one-to-one mapping and resource availability constraints. Concretely, virtual nodes in the same VN must be placed in different physical nodes and each physical node can host at most one virtual node. And the physical node $n_{p}$ must have enough available resources to place the corresponding virtual node $n_{v}$, i.e., $C(n_{v})\leq C(n_{p}),\forall n_{v}\in\mathcal{N}_{v}$. Link mapping aims to find a connected physical path $\rho_{p}=f_{\mathcal{L}}(l_{v})$ to route each virtual link $l_{v}$ while satisfying the path connectivity and resource capacity constraints. Concretely, the physical path $\rho_{p}$ should connect physical nodes hosting the two endpoints of virtual link $l_{v}$. And each physical link $l_{p}\in\rho_{p}$ of physical path $\rho_{p}$ should have enough bandwidth to route the corresponding virtual link $l_{v}$, i.e., $B(l_{v})\leq B(l_{p}),\forall l_{v}\in\mathcal{L}_{v},l_{p}\in\rho_{p}$.
Optimization Objective. Considering the randomness of online service requests, NFV-RA mainly aims to maximize the resource utilization of each instance $I$, which facilitates long-term resource profit and request acceptance (He et al., 2023; Zhang et al., 2023). To assess the solution quality $S=f_{\mathcal{G}}(I)$ of instance $I$, the revenue-to-cost ratio (R2C) is a widely used metric, defined as follows:
@@
Here, $\varkappa$ is a binary variable indicating the feasibility of the solution $S$: $\varkappa=0$ if $S$ violates some constraints, and $\varkappa=1$ otherwise. $\text{REV}(S)$ and $\text{COST}(S)$ denote the generated revenue and incurred cost by embedding VN $\mathcal{G}_{v}$. If $\varkappa=1$, $\text{REV}(S)$ denotes the revenue from the VN, calculated as $\sum_{n_{v}\in\mathcal{N}_{v}}{C}(n_{v})+\sum_{l_{v}\in\mathcal{L}_{v}}{B}(l_{v})$ and $\text{COST}(S)$ denotes the resource consumption of PN, calculated as $\sum_{n_{v}\in\mathcal{N}_{v}}{C}(n_{v})+\sum_{l_{v}\in\mathcal{L}_{v}}(|f_{\mathcal{L}}(l_{v})|\times{B}(l_{v}))$. Here, $|f_{\mathcal{L}}(l_{v})|$ quantifies the length of the physical path $\rho_{p}$ routing the virtual link $l_{v}$. See Appendix A.1 for the detailed formulation.
@@ 3 Virne: A Comprehensive NFV-RA Benchmark
To provide high-quality simulation and implementation, we introduce three design principles for Virne, following established software engineering practices (Van Vliet et al., 2008): (a) Versatile customization allows the platform to meet diverse simulation needs on varying network scenarios and conditions. (b) Scalable modularity built platforms with a modular architecture to support flexible configurations and easy extensibility. (c) Intuitive usability prioritizes a user-friendly interface, enabling users to focus on experiment outcomes rather than implementation complexities. Guided by these principles, as illustrated in Figure 2, we implement Virne with five key modules as follows.
@@ 3.1 Simulation Configuration
Modern network systems are diverse and intricate, associated with various realistic factors such as resource availability and service requirements. To support high customization for simulating different network scenarios and conditions, Virne abstracts both the PN and VN as graphs, with customizable attributes for nodes, links, and the overall network. Specifically, Virne provides the following key customizable elements: (a) Network Topologies. Users can select from various topological synthesis methods or real-world physical infrastructure topologies, such as those from SDNLib. (b) Resource Availability. Virne enables users to define multiple resource types (e.g., CPU, GPU, bandwidth) and their availability across different levels of the network (i.e., nodes, links and graph), allowing the reflection of the specific resource characteristics of different network environments (c) Service Requirements. Users can specify additional service requirements, such as latency, energy efficiency, and reliability, to ensure that the simulation reflects the needs of different network applications. By adjusting these parameters, Virne provides the flexibility to simulate a wide array of network scenarios, including cloud-based infrastructures, edge computing and 5G networks. This customization enables comprehensive testing, accommodating fluctuating resource demands and evolving network conditions. We provide configuration examples in Figure 6 and Appendix C.1.
@@ 3.3.1 NFV-RA as a Markov Decision Process
To address the randomness of online networks, most existing works model NFV-RA solution construction as a Markov Decision Process (MDP), which sequentially selects a physical node to place a virtual node at each decision step $t$, until all virtual nodes are placed or any constraints are violated. Decision sequence defaults to VN node index, with support for customizable node ranking in Virne.
@@
At each decision timestep $t$, observing the state $s_{t}$ of the environment, the agent takes an action $a_{t}\sim\pi(\cdot|s_{t})$ according to a policy $\pi_{\theta}$ parameterized by $\theta$. Then, the environment will transit to a new state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$ and feedback a reward $R(s_{t},a_{t})$ and. During interactions, a trajectory memory collects these step information as experiences. Using these experiences, the objective of the agent is to learn an optimal policy for resource allocation that maximizes the expected sum of discounted reward, i.e., the solution quality of NFV- $\pi_{\theta}^{*}=\arg\max_{\pi_{\theta}}\mathbb{E}_{\pi_{\theta}}\left[\sum_{i=0}^{T}\gamma^{i}R(s_{t},a_{t})\right]$
Note that the above MDP introduced is the most adopted version. As a holistic framework, Virne also supports the emerging MDP version in recent studies. Due to the presentation clarity and page limit, we offer other MDP versions of NFV-RA in Appendix C.2, such as Constrained MDP (Wang et al., 2025b) for constraint handling and Multi-task MDP (Wang et al., 2024c) for generalization.